OpenStudy (anonymous):

Derivative of P?? P(t)= A / 1+Be^(-kt). A, B, k are positive constants.

OpenStudy (anonymous):

Is this your function?\[P(t)=\frac{A}{1+Be^{-kt}}\]

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

first, re-write it as:\[P(t)=A(1+Be^{-kt})^{-1}\]

OpenStudy (anonymous):

Now let u=1+Be^(-kt) so du=-Bke^(-kt)dt...ok so far?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Now we have \[P(t)=Au^{-1}\]So taking derivative, using the chain rule, we get:\[dP/dt=(dP/du)(du/dt)\]or,\[dP/dt=-Au^{-2}(-Bke^{-kt})\]Getting rid of our new variable "u" we get:\[dP/dt=-A(1+Be^{-kt})^{-2}(-Bke^{-kt})\]I'll leave it for you to simplify if you wish.

OpenStudy (anonymous):

sorry, first line shoud really say P(u)=Au^(-1) as it is a function of u now, not t

OpenStudy (anonymous):

gotta run my little ones to daycare...I'll check back with you in a bit if you have questions on this one

OpenStudy (anonymous):

okayyy thanks!

OpenStudy (anonymous):

okay i got it it makes a lot more sense now.

OpenStudy (anonymous):

would you know how to evaluate the limit of P(t) toward infinity and -infinity?

OpenStudy (anonymous):

plug it in for t but I dont know what that gets me? 1+Be^-kinfinity?

OpenStudy (anonymous):

when it approaches infinity I got A/1 or A which is correct. However I keep getting it wrong when it approaches - infinity